OpenStudy (anonymous):

If the graph of f(x)=x^2+bx+c has its vertex at (-3,-11), find f(x). You have to find the quadratic equation.

OpenStudy (anonymous):

\[f(x)=x^2+bx+c\]

OpenStudy (anonymous):

the answer is: \[f(x)=x^2+6x-2\] but how do you get to the answer?????

OpenStudy (anonymous):

I got f(x)= x^2 + 6x - 8

OpenStudy (anonymous):

I already got it so long ago but thanx everyone!!!